• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 23:56:47 on April 9, 1999

    Entry number 17545

    keyword=Beam Time Accounting
    Hall A

    4/9/1999 1600-2400
    Experiment #: E91-010
    Shift Leader: Krishna Kumar
    Shift Workers: Richard Wilson (TO), Christian Cavati

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    45 60 10 55 0 25 33 60 4.8
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    15 0 50 5 60 35 27 0 3.2
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    60 60 60 60 60 60 50 60 7.83333333333333
    Planned Config-
    uration Change
    0 0 0 0 0 0 10 0 0.166666666666667
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 4 pm Laser intensity problem seems to be solved. Matt P. did something to the feedback between the laser amplifier. 7 pm. During system check the tune shift was changed to enable coil pulsing to work. The new tune shift is fine for coil pulsing but may be producing target boiling.

    Data Entry by: Richard Wilson